# Physics-Conditioned Diffusion Models for Lattice Gauge Theory

### v1

@@ 4 Physics-Conditioned Sampler
Sampling in diffusion models can be realized Wang:2023exq as a variant of stochastic quantization Parisi:1980ys , an approach for quantizing field theories that is particularly useful for gauge theories  Damgaard:1987rr ; Namiki:1993fd . In stochastic quantization, one introduces a fictitious time $\tau$, extending a field $\phi(x)$ to $\phi(x,\tau)$, which then evolves according to a Langevin equation,
@@
starting from a Gaussian prior for $\phi(x)$. By gradually decreasing $t$ from $T=1$ to 0 (where the noise vanishes), the resulting sample distribution becomes consistent with the target distribution used for training.
@@ 4.1 Physics-Conditioned Score Function
In this study, our target is to sample configurations at different values of the coupling $\beta$ using a diffusion model trained at a single specific coupling $\beta_{0}$, without requiring additional training. To achieve this, we construct a physics-conditioned sampler by drawing upon principles from statistical field theory. Configurations are sampled from the physical distribution, $p(\phi)=\exp[-S(\phi;\beta)]/Z$, where $\beta$ is a parameter (the inverse coupling in our case). If the action is separable as $S(\phi;\beta)=\beta\tilde{S}(\phi)$, one may perform simulations at a reference value $\beta_{0}$ and then use Ferrenberg-Swendsen or histogram reweighting PhysRevLett.61.2635 to obtain results for other values of $\beta$, provided the sampled ensembles have sufficient overlap. This idea has been proposed and validated for other field systems with different generative models Wang:2020nnsm ; Xu:2024tjp .
In diffusion models, the network-based score function $s_{\theta}$ is trained to approximate the drift term $\nabla_{\phi}\log p_{t}(\phi)$ in the reverse diffusion process. In contrast, in Langevin simulations of fields, the drift term is given by the functional derivative of the action, $-{\delta S[\phi]}/{\delta\phi(x,\tau)}$. If the diffusion model is well-trained, when the noise scale reaches 00, the score function should approximate the actual drift term, $s_{\hat{\theta}}(\phi(x,0),0;\beta)\simeq-{\delta S(\phi;\beta)}/{\delta\phi(x)}$. For separable actions, such as in pure gauge theories, both the action and its derivative scale linearly with the coupling $\beta$. This suggests that a well-trained score network at a reference coupling $\beta_{0}$ can directly replace the drift term in Langevin simulations at arbitrary $\beta$ when equilibrium is reached around the final steps. Then one can generate configurations at different values of $\beta$ from the trained diffusion model directly, using the sampler to be introduced in the next section.
@@ 4.2 Metropolis-Adjusted Annealed Langevin Algorithm
As explained earlier, generating samples for U(1) lattice gauge theory at different coupling values is only feasible if equilibrium is achieved in the final noise-scale steps. To ensure this, we use the annealed Langevin scheme proposed in Ref. song2019generative . In Fig. 4, the reverse denoising process is demonstrated along the horizontal time axis, with different noise-scales at different $t$. Each orange double arrow can be expanded into a dashed orange box describing the annealed scheme. This is where we run evolution along another time direction, $0<\tau\leq T_{A}$, with the fixed drift term $-\frac{\beta}{\beta_{0}}g^{2}(t)s_{\hat{\theta}}(\phi_{t},t)$. The final state of the annealed Langevin evolution is copied as the initial state at the next noise scale. By performing multiple Langevin updates at each progressively decreasing noise scale $t_{i}$, one ensures that the distribution approaches $p(\phi,t_{i})\propto e^{-S_{DM}(\phi,t_{i})}$, where $\partial S_{DM}(\phi,t_{i})/\partial\phi=-\frac{\beta}{\beta_{0}}\hat{s}_{\theta}(\phi,t_{i})$. This will eventually bring us to $S_{DM}(\phi,t_{i}\rightarrow 0)\simeq S(\phi;\beta)$ in the final steps. The detailed set-up of the annealed scheme can be found in Appendix A.
@@
[…]
In a general MALA roberts1998optimal ; girolami2011riemann , from the current state $\phi_{\tau}$, the proposed update is
@@
where $\alpha_{i}$ is a redefined noise-scale parameter which can be found in Appendix A. The accept-reject step is
@@
where $p(\phi)$ is computed from the known physical distribution. Note that conditioned on $\phi_{\tau}$, the proposal step boils down to drawing a Gaussian with mean $\phi_{\tau}+\alpha_{i}f(\phi_{\tau},\tau)$ and variance $\sqrt{2\alpha_{i}}I_{\tau}$. Hence, the proposal kernel has the explicit form
@@
where $n$ is the number of degrees of freedom, which is $n=V$ in our case. The transition probability can then be computed explicitly, as
@@
Combined with the annealed scheme, the Metropolis-adjusted annealed Langevin algorithm can generate exact configurations at different inverse couplings $\beta$. A detailed set-up can be found in Algorithm 2 of Appendix A. In fact, the classical Langevin simulation is obtained by replacing the time-dependent score function by the physical drift term derived from the given action. We will compare our results with the Langevin simulation according to Eq. (18). In this study, the results of the diffusion model are all generated by the Metropolis-adjusted Langevin annealed algorithm.
@@ 5 Numerical Experiments
In this work, configurations for training are generated by HMC, which is an efficient method for generating gauge field configurations at least at $\beta=1.0$. To generate better samples, for each Markov chain in HMC, we have a “200-step burn-in loop” for thermalization, and after 128 additional steps, we select the final step of the 128-step update as a configuration. Here one step indicates one MC trajectory with a well-established HMC integrator, as in Ref. Wang:2023exq .
@@ 5.1.1 Training and validation
First, we evaluate the training performance in the topologically unfrozen region at $\beta=1$. All configurations are generated by the original sampler, as introduced in Ref. Wang:2023exq , from the trained diffusion model. In Fig. 5, the distributions of the Wilson loop (left) and the topological charge (right) are shown as orange bins, demonstrating consistency with the HMC results, which are represented by blue envelopes and bins. Qualitatively, the diffusion model achieves comparable performance to HMC in estimating both local and topological observables in the topologically unfrozen region. More quantitative results are shown in the following sections with our new sampler.
@@ 5.1.2 Generation at different lattice sizes
HMC DM Langevin Exact HMC DM Langevin Exact
@@ 5.2.1 Mitigate freezing
Loop size ($l$) HMC DM Langevin Exact
@@
Using the same trained diffusion model, we generate 1024 configurations at $\beta=7$ and compare them with a set of 1024 configurations generated via HMC. The results are presented in Fig. 6. We observe that the Wilson loop distributions (left) remain consistent between the two approaches. A more detailed comparison of Wilson loop at different sizes is provided in Table 2, which also includes a comparison with straightforward Langevin simulations. However, for the topological charge in Fig. 6 (right), HMC exhibits topological freezing, sampling only values of the topological charge $Q$ near or at zero. In contrast, the physics-conditioned sampler using the DM explores a broader range of topological sectors, leading to a larger topological susceptibility that is consistent with the analytical results. In addition, one can also estimate the likelihood for different conditions to evaluate whether the model is approaching the target physics distribution, which is discussed in Appendix D.
@@ 5.2.2 Generation at different lattice sizes and couplings
Next we estimate observables at both different couplings and lattice sizes. Table 3 presents a comparison of observables at fixed coupling $\beta=7$ for lattice sizes $L=8\times 8,16\times 16$ and $32\times 32$. The observables studied are the $1\times 1$ Wilson loop and the topological susceptibility. For the three samplers, the Wilson loop values remain relatively stable across different lattice sizes and numerical methods, with all deviations within statistical uncertainties. However, the topological susceptibility shows significant variation depending on the method used. In particular, Langevin dynamics yields notably higher susceptibility values compared to the exact results, while the HMC results indicate freezing. The DM remains closer to the exact predictions.
@@
HMC DM Langevin Exact HMC DM Langevin Exact
8 0.927(13) 0.926(13) 0.921(13) 0.926 0.00006(3) 0.0040(12) 0.0143(5) 0.0040
16 0.926(7) 0.926(7) 0.924(6) 0.926 0.00013(2) 0.0043(5) 0.0131(5) 0.0039
32 0.926(3) 0.925(4) 0.924(4) 0.926 0.00013(2) 0.0040(4) 0.0137(6) 0.0039
@@
HMC DM Langevin Exact HMC DM Langevin Exact
@@
While the diffusion model maintains close agreement with the exact values, the HMC results exhibit again topological freezing, and the Langevin method systematically overestimates the susceptibility, particularly at higher $\beta$. These findings underscore the effectiveness of physics-conditioned diffusion models in computing lattice gauge theory observables.
@@ 6 Summary and Outlook
In this study, we applied diffusion models to generate gauge field configurations. The correspondence between the stochastic differential equations encountered in diffusion models and in stochastic quantization enables the model to generate configurations at different lattice sizes and coupling constants without additional retraining. To ensure that the sampled distribution corresponds to the physical distribution, we combined the Metropolis-adjusted Langevin algorithm with the annealed Langevin sampler. Our results demonstrate that this multi-chain sampler effectively mitigates topological freezing at large $\beta$, and that the diffusion model proposes improved configurations for the Markov chains, leading to estimates of observables that remain consistent with exact results.
In our study, the diffusion model performs well for Abelian pure gauge fields, even with a simple U-net architecture. Extending this approach to non-Abelian gauge theories, including fermionic systems, appears straightforward; however, the non-locality of fermion determinants remains a significant challenge. The first attempt for a complex action (sign) problem has been reported recently Habibi:2024fbn . The integration of diffusion models with neural networks specifically designed for lattice configurations Favoni:2020reg ; Nagai:2021bhh ; Nagai:2025rok ; Aarts:2025gyp is also a promising direction. The scalability of diffusion models suggests its potential for generating configurations on larger lattices at large $\beta$, requiring training only on samples from smaller lattice sizes and lower $\beta$ values. Our objective for the future is to develop a complementary approach that enhances the efficiency and accuracy of lattice Quantum Chromodynamics computations, particularly in regimes where traditional techniques face limitations.
@@ Acknowledgements.
We thank Shiyang Chen and Diaa E. Habibi for comments and Drs. Akinori Tanaka, Tetsuo Hatsuda, Andreas Ipp, Gurtej Kanwar, Fernando Romero-López, and Shuzhe Shi for helpful discussions. We thank the DEEP-IN working group at RIKEN-iTHEMS for support in the preparation of this paper. WW and QZ are supported in part by Natural Science Foundation of China under grants No. 12125503 and 12335003, and SJTU Kunpeng & Ascend Center of Excellence. GA is supported by STFC Consolidated Grant ST/X000648/1. KZ is supported by the CUHK-Shenzhen university development fund under grant No. UDF01003041 and UDF03003041, and Shenzhen Peacock fund under No. 2023TC0179. LW is also supported by the RIKEN TRIP initiative (RIKEN Quantum) and JST-BOOST grant. Research Data and Code Access – The code and data used for this manuscript can be found at https://github.com/zzzqt/DM4U1. Open Access Statement – For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.

### v2

@@ 4 Physics-Conditioned Sampler
Sampling in diffusion models can be realized Wang et al. (2024) as a variant of stochastic quantization Parisi et al. (1981), an approach for quantizing field theories that is particularly useful for gauge theories  Damgaard and Huffel (1987); Namiki (1993). In stochastic quantization, one introduces a fictitious time $\tau$, extending a field $\phi(x)$ to $\phi(x,\tau)$, which then evolves according to a Langevin equation,
@@
starting from a Gaussian prior for $\phi(x)$. By gradually decreasing $t$ from $T=1$ to 0 (where the noise vanishes), the resulting sample distribution controlled by the score function becomes consistent with increasingly close to the target distribution used for training.
@@ 4.1 Physics-Conditioned Score Function
In this study, our target is to sample configurations at different values of the coupling $\beta$ using a diffusion model trained at a single specific coupling $\beta_{0}$, without requiring additional training. To achieve this, we construct a physics-conditioned sampler by drawing upon principles from statistical field theory. Configurations are sampled from the physical distribution, $p(\phi)=\exp[-S(\phi;\beta)]/Z$, where $\beta$ is a parameter (the inverse coupling in our case). In the pure gauge case, the action is separable as $S(\phi;\beta)=\beta\tilde{S}(\phi)$, and one can easily perform simulations at a reference value $\beta_{0}$ and then use Ferrenberg-Swendsen or histogram reweighting Ferrenberg and Swendsen (1988) to obtain results for other values of $\beta$, provided the sampled ensembles have sufficient overlap. This idea has been proposed and validated for other field systems with different generative models Wang et al. (2020); Xu et al. (2024).
In diffusion models, the network-based score function $s_{\theta}$ is trained to approximate the drift term $\nabla_{\phi}\log p_{t}(\phi)$ in the reverse diffusion process. In contrast, in Langevin simulations of fields, the drift term is given by the functional derivative of the action, ${\delta S[\phi]}/{\delta\phi(x,\tau)}$. If the diffusion model is well-trained, when the noise scale reaches 0, the score function should approximate the actual drift term, $s_{\hat{\theta}}(\phi(x,0),0;\beta)\simeq-{\delta S(\phi;\beta)}/{\delta\phi(x)}$. For separable actions, such as in pure gauge theories, both the action and its derivative scale linearly with the coupling $\beta$. This suggests that a well-trained score network at a reference coupling $\beta_{0}$ can directly replace the drift term in Langevin simulations at arbitrary $\beta$ when equilibrium is reached around the final steps. Then one can generate configurations at different values of $\beta$ from the trained diffusion model directly, using the sampler to be introduced in the next section.
@@ 4.2 Metropolis-Adjusted Annealed Langevin Algorithm
As explained earlier, generating samples for U(1) lattice gauge theory at different coupling values is only feasible if equilibrium is achieved in the final noise-scale steps. To ensure this, we use the annealed Langevin scheme proposed in Ref. Song and Ermon (2019). In Fig. 4, the reverse denoising process is demonstrated along the horizontal time axis, with different noise-scales at different $t$. Each orange double arrow can be expanded into a dashed orange box describing the annealed scheme. This is where we run evolution along another time direction, $0<\tau\leq T_{A}$, with the fixed drift term $\frac{\beta}{\beta_{0}}g^{2}(t)s_{\hat{\theta}}(\phi_{t},t)$. The final state of the annealed Langevin evolution is copied as the initial state at the next noise scale. By performing multiple Langevin updates at each progressively decreasing noise scale $t_{i}$, one ensures that the distribution approaches $p(\phi,t_{i})\propto e^{-S_{DM}(\phi,t_{i})}$, where $\partial S_{DM}(\phi,t_{i})/\partial\phi=-\frac{\beta}{\beta_{0}}\hat{s}_{\theta}(\phi,t_{i})$. This will eventually bring us to $S_{DM}(\phi,t_{i}\rightarrow 0)\simeq S(\phi;\beta)$ in the final steps. The detailed set-up of the annealed scheme can be found in Appendix A.
@@
In addition, to ensure that a Markov chain has the correct stationary measure, one can add a Metropolis-Hastings filter, i.e., an accept-reject step, along it. We extended the Metropolis-adjusted Langevin algorithm (MALA) Roberts and Rosenthal (1998); Radivojević and Akhmatskaya (2020) to Metropolis-Adjusted Annealed Langevin algorithm (MAALA) by building multi-Markov chains with the annealed process to update configurations and adjusting the chains by Metropolis updates during the final steps, $0\leq t\leq 0.001$. [^1] In our case, $t\leq 0.001$ means $N_{C}=35$ discrete time steps along the diffusion model backward direction in the set-up of noising scheme, using $0.9^{N_{T}-N_{C}}\leq 0.001$. More details can be found in Appendix A. This will prepare $N_{C}$ chains for collecting configurations to estimate observables accurately. This sampler will be referred as DM-MAALA (DM for short in the following context). As shown in the dashed blue box of Fig. 4, the Metropolis-adjusted procedure is employed in the annealed process for the last $N_{C}$ noise-scales. Since $S_{DM}(\phi,t_{i}\rightarrow 0)\simeq S(\phi;\beta)$, the sampler can reach a reasonably high acceptance rate as shown in Appendix B.
In a general MALA Roberts and Rosenthal (1998); Radivojević and Akhmatskaya (2020), from the current state $\phi_{\tau}$, the proposed update is
@@
where $\alpha_{i}$ is a redefined noise-scale parameter which can be found in Appendix A, and $F$ is one proposal kernel, which is the score function in DM. The accept-reject step is
@@
where $p(\phi)$ is computed from the known physical distribution. Note that conditioned on $\phi_{\tau}$, the proposal step boils down to drawing a Gaussian with mean $\phi_{\tau}+\alpha_{i}F(\phi_{\tau},\tau)$ and variance $2\alpha_{i}I_{\tau}$. Hence, the proposal kernel has the explicit form
@@
where $n$ is the number of degrees of freedom, which is $n=2V$ in our case. The transition probability can then be computed explicitly, as
@@
Combined with the annealed scheme, MAALA can generate exact configurations at different inverse couplings $\beta$. A detailed set-up can be found in Algorithm 2 of Appendix A. In fact, the classical MALA simulation is obtained by replacing the time-dependent score function by the time-independent physical drift term derived from the given action. We will compare our results with the MALA simulation according to Eq. (19). In this study, the results of the diffusion model are all generated by the Metropolis-adjusted Langevin annealed algorithm.
@@ 5 Numerical Experiments
In this work, configurations for training are generated by HMC on one single chain , which is an efficient method for generating gauge field configurations, at least at $\beta=1.0$. To generate better samples, for each Markov chain in HMC, we have a “200-step burn-in” steps222“Burn-in” in this study refers to the generation steps before Metropolis is employed. for thermalization, and after 128 additional steps, we select the final step of the 128-step update as a configuration. Here one step indicates one MC trajectory with a well-established HMC integrator, as in Ref. Wang et al. (2024).
A comprehensive comparison between the algorithms discussed here can be found in Appendix E.
@@ 5.1.1 Training and validation
First, we evaluate the training performance in the topologically unfrozen region at $\beta=1$. All configurations are generated by the original sampler, as introduced in Ref. Wang et al. (2024), from the trained diffusion model. In Fig. 5, the distributions of the Wilson loop (left) and the topological charge (right) are shown as orange bins, demonstrating consistency with the HMC results, which are represented by blue envelopes and bins. Qualitatively, the diffusion model achieves comparable performance to HMC in estimating both local and topological observables in the topologically unfrozen region. More quantitative results are shown in the following sections with our new sampler.
@@ 5.1.2 Generation at different lattice sizes
HMC DM MALA Exact HMC DM MALA Exact
@@ 5.2.1 Mitigate freezing
Loop size ($l$) HMC DM MALA Exact
@@
Using the same trained diffusion model, we generate 1024 configurations at $\beta=7$ and compare them with a set of 1024 configurations generated via HMC. The results are presented in Fig. 6. We observe that the Wilson loop distributions (left) remain consistent between the two approaches. A more detailed comparison of Wilson loop at different sizes is provided in Table 2, which also includes a comparison with MALA.
However, for the topological charge in Fig. 6 (right), HMC exhibits topological freezing, sampling only values of the topological charge $Q$ near or at zero – see Appendix E for a comparison between all the algorithms discussed above. In contrast, the physics-conditioned sampler using the DM explores a broader range of topological sectors, leading to a larger topological susceptibility that is consistent with the analytical results. In addition, one can also estimate the likelihood for different conditions to evaluate whether the model is approaching the target physics distribution, which is discussed in Appendix D.
@@ 5.2.2 Generation at different lattice sizes and couplings
Next we estimate observables at both different couplings and lattice sizes. Table 3 presents a comparison of observables at fixed coupling $\beta=7$ for lattice sizes $L=8\times 8,16\times 16$ and $32\times 32$. The observables studied are the $1\times 1$ Wilson loop and the topological susceptibility. For the three samplers, the Wilson loop values remain relatively stable across different lattice sizes and numerical methods, with all deviations within statistical uncertainties. However, the topological susceptibility shows significant variation depending on the method used. In particular, MALA yields notably higher susceptibility values compared to the exact results, while the HMC results indicate freezing. The DM remains closer to the exact predictions.
@@
HMC DM MALA Exact HMC DM MALA Exact
8 0.927(13) 0.926(13) 0.921(13) 0.926 0.00006(3) 0.0057(2) 0.0143(5) 0.0040
16 0.926(7) 0.926(7) 0.924(6) 0.926 0.00013(2) 0.0063(3) 0.0131(5) 0.0039
32 0.926(3) 0.925(4) 0.924(4) 0.926 0.00013(2) 0.0051(3) 0.0137(6) 0.0039
@@
HMC DM MALA Exact HMC DM MALA Exact
@@
We found that for the $1\times 1$ Wilson loop, a DM trained at $\beta=1$ yields results consistent with the exact values when employed at larger beta values (we went up to $\beta=11$), underscoring the effectiveness of physics-conditioned diffusion models. For the topological susceptibility on the other hand, such an extrapolation was found to break down after $\beta\gtrsim 5$, leaving a more limited range of couplings available. Despite best efforts, this issue is currently an open problem.
@@ 6 Summary and Outlook
In this study, we applied diffusion models to generate gauge field configurations. The correspondence between the stochastic differential equations encountered in diffusion models and in stochastic quantization enables the model to generate configurations at different lattice sizes and coupling constants without additional retraining. To ensure that the sampled distribution corresponds to the physical distribution, we combined the Metropolis-adjusted Langevin algorithm with the annealed Langevin sampler. Our results demonstrate that this multi-chain sampler effectively mitigates topological freezing at large $\beta$, and that the diffusion model proposes improved configurations for the Markov chains, leading to estimates of observables that remains consistent withe exact results when extrapolation is in intermediate interval.
Although DM-MAALA shows good ergodicity across topological sectors in early sampling stages Zhu et al. (2024), the DM generation cannot completely avoid the bias of topology distribution in the training set. Intuitively speaking, when extrapolation to larger $\beta$ values, only the magnitude of the learned score is changed, while the direction is unaffected. This may cause the generated configurations to remain biased towards the topology distribution of the training set. Going forward, it is crucial to investigate this further (and resolve it). To mitigate this, one possibility is to adopt topological‐sector–fixed sampling schemes Fukaya (2006) as current applications in HMC methods, and later merge or stitch configurations across sectors. Another practical improvement is to restrict extrapolation to small $\beta$ intervals so that sectors overlap sufficiently and the sampler can jump among them easily in the Metropolis procedure. Furthermore, it is also possible to introduce auxiliary fields or coupling variables to sample $Q$ directly Chen (2024). Although it will introduce additional cost to the sampling procedure, it may offer a promising solution in our framework.
Apart from this difficulty, extending this approach to non-Abelian gauge theories, including fermionic systems, appears straightforward; however, the non-locality of fermion determinants in QCD-like theories remains a significant challenge as it may need a novel structure of network architecture to catch long-range correlations. The first attempt for a complex action (sign) problem has been reported recently Habibi et al. (2025). For sampling, improvements of annealed Langevin dynamics Albergo and Vanden-Eijnden (2024) should be considered. Concerning the neural network architecture, the integration of diffusion models with neural networks specifically designed for lattice configurations Favoni et al. (2022); Nagai and Tomiya (2021); Nagai et al. (2025); Aarts et al. (2025a) is a promising direction. The scalability of diffusion models suggests its potential for generating configurations on larger lattices at large $\beta$, requiring training only on samples from smaller lattice sizes and lower $\beta$ values. Our objective for the future is to develop a complementary approach that enhances the efficiency and accuracy of lattice Quantum Chromodynamics computations, particularly in regimes where traditional techniques face limitations.
Notes – A related study has been released recently Ranner (2024).
@@ Acknowledgements.
We thank Shiyang Chen and Diaa E. Habibi for comments and Drs. Akinori Tanaka, Tetsuo Hatsuda, Andreas Ipp, Gurtej Kanwar, Fernando Romero-López, and Shuzhe Shi for helpful discussions. We thank the DEEP-IN working group at RIKEN-iTHEMS for support in the preparation of this paper. WW and QZ are supported in part by Natural Science Foundation of China under grants No. 12125503 and 12335003, and SJTU Kunpeng & Ascend Center of Excellence. GA is supported by STFC Consolidated Grant ST/X000648/1. KZ is supported by the CUHK-Shenzhen university development fund under grant No. UDF01003041 and UDF03003041, Ministry of Science and Technology of China under Grant No. 2024YFA1611004, NSFC fund under No. 92570117 and Shenzhen Peacock fund under No. 2023TC0179. L. Wang is also supported by the RIKEN-TRIP initiative (RIKEN-Quantum), JSPS KAKENHI Grant No. 25H01560, and JST-BOOST Grant No.JPMJBY24H9. Research Data and Code Access – The code and data used for this manuscript can be found at https://github.com/zzzqt/DM4U1. Open Access Statement – For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.